OpenStudy (anonymous):

factor the given polynominal PLEASE HELP 15x^2+7x-4

OpenStudy (anonymous):

(3x-1)(5x+4)

OpenStudy (anonymous):

thank u so mch. can u help with one more? may i give it to you?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

find the zeros of f(x)=x^2+6x+3

OpenStudy (anonymous):

oki so make it equal to zero and solve

OpenStudy (anonymous):

i dont know how find x when there are two x's

OpenStudy (anonymous):

x(x+6)+3

OpenStudy (anonymous):

got it.

OpenStudy (anonymous):

thats a triky one

OpenStudy (anonymous):

so now what?

OpenStudy (anonymous):

u gotta find an x

OpenStudy (anonymous):

-9?

OpenStudy (anonymous):

\[\sqrt{6}-3 and -3-\sqrt{{6}}\]

OpenStudy (anonymous):

its kind a weird one

OpenStudy (anonymous):

thank you so much. u really helped. I have an easier one if you would likfe to help again... :)

OpenStudy (anonymous):

do you know the quadratic formula? It works when "normal" factoring doesn't seem easy.